OpenStudy (anonymous):

find the sum in simplest form. 6m^2/18mn + 3n^2/18mn

OpenStudy (anonymous):

\[\frac{ 6m^2 }{ 18mn }+\frac{ 3n^2 }{ 18mn }\]

Directrix (directrix):

These are two algebraic fractions with the same denominator. So, add the numerators..

OpenStudy (anonymous):

\[\frac{ 9m^2n^2 }{ 18mn }\]

Directrix (directrix):

No, the 6m^2 + 3n^2 are not like terms and cannot be added.

Directrix (directrix):

What is in the attachment is not the final answer.

Directrix (directrix):

Look at this

Directrix (directrix):

That is not the final answer because the 3 in the numerator and the 18 in the denominator have a common factor of 3.

OpenStudy (anonymous):

\[\frac{ 2m^2+n^2 }{ 6mn }\]

Directrix (directrix):

That is what I got.

Directrix (directrix):

I guess I'll have to learn LaTex. Your posts look professional next to my handwriting.

OpenStudy (anonymous):

im sorry iv been taught from the begining everything has to be in order and neatly but i cant spell

Directrix (directrix):

I can spell but the lessons on handwriting did not faze me. You should be proud, not sorry. I was thinking that if I try to learn LaTex, I can ask you for help.

OpenStudy (anonymous):

thank you, you see the blue equation button next to draw?

Directrix (directrix):

\[2\div4=\sqrt{1/4}\]

OpenStudy (anonymous):

exclaty

Directrix (directrix):

I'll have to try it in the next problem you post.

OpenStudy (anonymous):

okay ill post now

Directrix (directrix):

okay